### v1

@@ 3 Optimality Conditions
In this section, we introduce two important concepts of stationarity for problem (1), namely, the optimization stationarity and game stationarity, and examine their relationships. To this end, we first definite the value function $\Phi:\mathbb{E}\to\mathbb{R}$ by
@@
The next proposition concerns the value function $\Phi$. Recalled that, a function $\varphi:\mathbb{E}\to\mathbb{R}$ is said to be $\ell$-weakly convex if $\varphi(\cdot)+(\ell/2)\|\cdot\|^{2}$ is convex.
@@
Let $y_{x}\in\arg\max_{y\in S}\{f(x,y)-g(y)\}$. The value function $\Phi$ is $L_{x}$-weakly convex and for all $x\in\mathcal{M}$, it holds that $\nabla_{x}f(x,y_{x})\in\partial\Phi(x)$. Moreover, if $f(x,y)-g(y)$ is strongly concave in $y$, then $\Phi$ is differentiable and $\nabla\Phi(x)=\nabla_{x}f(x,y_{x})$.
@@
Because $f(x,y)$ is $L_{x}$-smooth respect to $x$, $f(x,y)+\frac{L_{x}}{2}\left\|x\right\|^{2}-g(y)$ is convex in $x$ for any given $y\in S$. Thus, $\Phi(x)+\frac{L_{x}}{2}\left\|x\right\|^{2}=\max_{y\in S}f(x,y)+\frac{L_{x}}{2}\left\|x\right\|^{2}-g(y)$ is convex. Therefore, $\Phi(x)$ is $L_{x}$-weakly convex. Since $S$ is compact, according to Danskin’s theorem (bernhardTheoremDanskinApplication1995, , Theorem D1), $\nabla_{x}f(x,y_{x})\in\partial\Phi(x)$, and in particular, $\nabla\Phi(x)=\nabla_{x}f(x,y_{x})$ when $f(x,y)-g(y)$ is strongly concave in $y$. ∎
In view of Proposition 2 and (1983Optimization, , Proposition 2.4.3), we see that any local minimizer $x^{*}$ of problem (9) must satisfy
@@
This naturally leads to the following definition of optimization stationarity.
@@
    Let $\varepsilon>0$ be given. Inspired by perez2019subdifferential , we say that $x^{*}$ is an $\varepsilon$-optimization-stationary point of problem (1) if
    $\mathrm{dist}(0,\partial_{\varepsilon}(\Phi+h)(x^{*})+\partial\delta_{\mathcal{M}}(x^{*}))<\varepsilon,$
    with
    $\partial_{\varepsilon}(\Phi+h)(x^{*}):=\{z\in\mathbb{E}\mid\langle z,v\rangle-\varepsilon\leq(\Phi+h)^{\circ}(x^{*};v),~{}\forall v\in\mathbb{E}\}.$
@@
Suppose that $x^{*}$ is an optimization-stationary point of problem (1). By Proposition (2), $\Phi(x)+h(x)+L_{x}\|x-x^{*}\|^{2}$ is strongly convex. As $0\in\partial\delta_{\mathcal{M}}(x^{*})+\partial\Phi(x^{*})+\partial h(x^{*})$ and $\partial\delta_{\mathcal{M}}(x^{*})=N_{x^{*}}{\mathcal{M}}=\partial\delta_{T_{x^{*}}{\mathcal{M}}}(0)$, then $x^{*}$ is the unique minimum point of the problem
@@
$f(x^{*},y^{*})+h(x^{*})-g(y^{*})$ $=\inf_{\begin{subarray}{c}x\in A_{x^{*}}{\mathcal{M}},\\ \|x-x^{*}\|\leq 1\end{subarray}}\sup_{y\in S}f(x,y)+h(x)-g(y)+L_{x}\|x-x^{*}\|^{2}$
@@
Conversely, suppose that $(x^{*},y^{*})$ is a game stationary point of problem (1). Then, according to Proposition 2, $y^{*}=\arg\max_{y\in S}f(x^{*},y)-g(y)$ and $\nabla_{x}f(x^{*},y^{*})\in\partial\Phi(x^{*})$. Thus, $0\in\partial\Phi(x^{*})+\partial h(x^{*})+\partial\delta_{\mathcal{M}}(x^{*})$. ∎
@@
where $\beta>0$ and $u=\arg\min_{v\in T_{x}\mathcal{M}}\langle\nabla_{x}f(x,y),v\rangle+h(x+v)+\frac{\beta}{2}\|v\|^{2}$. This definition is motivated in part by the stationarity measure used in Euclidean minimax problem luHybridBlockSuccessive2020 . It is straightforward to verify that $(x^{*},y^{*})$ is a game-stationary point of problem (1) if and only if there exists $\beta>0$ such that $\mathcal{G}^{\beta}(x^{*},y^{*})=0$. Using this measure, we can also introduce the notion of $\varepsilon$-game-stationary point.
@@
Let $\varepsilon>0$ be given. Suppose that $(x^{*},y^{*})$ is an $\varepsilon$-game-stationary point of problem (1) with $\mathcal{G}^{\beta}(x^{*},y^{*})\leq\varepsilon$. Then $x^{*}$ is a $(C\varepsilon)$-optimization-stationary poin, where $C=(2\sigma_{y}^{2}+1)+2L_{y}\sigma_{y}/\beta$ and $\sigma_{y}:=\sup_{y\in S}\left\|y\right\|$.
@@
Suppose the point $(x^{*},y^{*})$ is an $\varepsilon$-game-stationary point of problem (1), then there exists $z$ such that $\|z\|=1$ and $\varepsilon z\in\nabla_{y}f(x^{*},y^{*})-\partial g(y^{*})-\partial\delta_{S}(y^{*})$. For $u=\arg\min_{v\in T_{x}\mathcal{M}}\langle\nabla_{x}f(x,y),v\rangle+h(x+v)+\frac{\beta}{2}\|v\|^{2}$, let
$\Phi_{\varepsilon}^{u}(x):=$ $\max_{y\in S}f(x+u,y)-g(y)-\frac{\varepsilon}{2}\|y-(y^{*}-z)\|^{2}$
$+\langle\nabla_{y}f(x^{*},y^{*})-\nabla_{y}f(x^{*}+u,y^{*}),y\rangle.$
By Proposition 2 and the choice of $z$, we have $\nabla\Phi_{\varepsilon}^{u}(x^{*})=\nabla_{x}f(x^{*}+u,y^{*})$ and
$\Phi_{\varepsilon}^{u}(x^{*})+\langle\nabla_{x}f(x^{*}+u,y),x-(x^{*}+u)\rangle\leq\Phi_{\varepsilon}^{u}(x-u)+\frac{L_{x}}{2}\|x-(x^{*}+u)\|^{2}.$
Because
$\Phi(x^{*}+u)-\varepsilon(2\sigma_{y}^{2}+1)-L_{y}\sigma_{y}\|u\|\leq\Phi_{\varepsilon}^{u}(x^{*})$
and
$\Phi_{\varepsilon}^{u}(x-u)\leq\Phi(x)+L_{y}\sigma_{y}\|u\|,$
we arrive at
$\Phi(x^{*}+u)+\langle\nabla_{x}f(x^{*}+u,y),x-(x^{*}+u)\rangle-\varepsilon(2\sigma_{y}^{2}+1)-2L_{y}\sigma_{y}\|u\|$
$\leq$ $\Phi(x)+\frac{L_{x}}{2}\|x-(x^{*}+u)\|^{2}.$
Since $h$ is convex and $L_{h}$-Lipschitz continuous, then
$h(x^{*}+u)+\langle\xi,x-(x^{*}+u)\rangle\leq h(x)$
for all $\xi\in\partial h(x^{*}+u)$. Therefore, by combing the above two inequalities, we obtain
$\nabla_{x}f(x^{*}+u,y^{*})+\partial h(x^{*}+u)\subset\partial_{\hat{\varepsilon}}(\Phi+h)(x^{*}+u),$
with $\hat{\varepsilon}=\varepsilon(2\sigma_{y}^{2}+1)+2L_{y}\sigma_{y}\|u\|$. Furthermore, as $\|\mathcal{G}_{x}^{\beta}(x^{*},y^{*})\|\leq\varepsilon$, we can conclude that $x^{*}$ is a $(C\varepsilon)$-optimization-stationary point with $C=(2\sigma_{y}^{2}+1)+2L_{y}\sigma_{y}/\beta$. ∎
4 The proposed MPGDA algorithm
In this section, we propose our MPGDA method for solving problem (1) and establish its iteration complexity. Our algorithm development is partially motivated by heApproximationProximalGradient2024 ; luHybridBlockSuccessive2020 ; xuRiemannianAlternatingDescent2024 . For the $k$-th iteration, we introduce the following value function:
@@ 4 The proposed MPGDA algorithm
where $\Phi_{k}:\mathbb{E}_{1}\to\mathbb{R}$ is defined at $x\in\mathbb{E}_{1}$ by
@@
Here, $y_{k}$ denotes the iterate with respect to $y$ obtained in the previous iteration, $\gamma_{k}>0$ is the regularization parameter, and $\rho_{k}>0$ is the proximal parameter. Moreover, we choose $\gamma_{k}=\gamma_{0}/k^{1/3}$ with some $\gamma_{0}>0$ and assume the non-increasing sequence $\{\rho_{k}\}$ to be summable, i.e.,
@@
We further introduced the mapping $\bar{y}_{k}:\mathbb{E}_{1}\to\mathbb{E}_{2}$ defined by
$\bar{y}_{k}(x):=\arg\max_{y\in S}\{f(x,y)-g(y)-\frac{\gamma_{k}}{2}\|y\|^{2}-\frac{\rho_{k}}{2}\|y-y_{k}\|^{2}\}.$ (15)
Then it holds that $Q_{k}(x)=F(x,\bar{y}_{k}(x))-\frac{\gamma_{k}}{2}\|\bar{y}_{k}(x)\|^{2}-\frac{\rho_{k}}{2}\|\bar{y}_{k}(x)-y_{k}\|^{2}$ and in view of Danskin’s theorem (danskinTheoryMaxMinApplications1966, , Theorem D1), $\Phi_{k}$ is smooth on $\mathbb{E}_{1}$ and $\nabla\Phi_{k}(x)=\nabla_{x}f(x,\bar{y}_{k}(x))$.
Now, at the $k$-th iteration, the proposed MPGDA method computes the new iterate by approximately solving the subproblem:
@@
Inspired by the structure of $Q_{k}$ and the ManPG algorithm (chenProximalGradientMethod2020, , Algorithm 4.1), we perform Riemannian manifold proximal gradient descent steps for $T_{k}$ times to generate $x_{k+1}$ with $T_{k}$ being an integer in $[1,\overline{T}]$ where $\overline{T}$ is a predetermined positive integer. Then $y_{k+1}$ is updated as $\bar{y}_{k}(x_{k+1})$ with $\bar{y}_{k}$ defined in (15). Notice that the update of $y$ can also be interpreted as performing a proximal ascent step on the function $f(x_{k+1},y)-g(y)-\frac{\gamma_{k}}{2}\|y\|^{2}$. We formally presented the proposed method in Algorithm 1.
Algorithm 1 Manifold Proximal Gradient Descent Ascent (MPGDA) for problem (1)
0:  $c_{1}>0$, $\eta\in(0,1)$, $\gamma_{0}>0$, $x_{0}\in\mathcal{M}$, $y_{0}\in\mathrm{dom}(g)\cap S$, $0<l_{\min}<l_{\max}$, $\{\rho_{k}\}\in\mathcal{S}$, $\{\gamma_{k}=\gamma_{0}/k^{1/3}:k\in N\}$, $\{T_{k}\}\subset[1,\overline{T}]$;
1:  for $k=0,\dots$ do
2:     Set $x_{k,0}=x_{k}$,
@@
$v_{k,i}:=\arg\min_{v\in T_{x_{k,i}}\mathcal{M}}\langle\nabla_{x}f(x_{k,i},\bar{y}_{k}(x_{k,i})),v\rangle+h(x_{k,i}+v)+\frac{\beta_{k,i}}{2}\left\|v\right\|^{2},$ (17)
@@
Before proceeding, we make two remarks on the iteration procedure of MPGDA algorithm. First, given $x_{k,i}\in\mathcal{M}$, MPGDA algorithm computes a descent direction $v_{k,i}$ of $Q_{k}$ restricted to the tangent space $T_{x_{k,i}}\mathcal{M}$ via tackling the linear constrained strongly convex optimization problem (17), which can usually be efficiently solved, see (li2024proximal, , Section 4), for example. Second, in practice, we make some choices of parameters $l_{k,i}$ and $\rho_{k}$, which enhance the efficiency of MPGDA algorithm in our test and also satisfy $l_{k,i}\in[l_{\min},l_{\max}]$ and $\rho_{k}\in S$. For parameters $l_{k,i}$, we adopt the Riemannian Barzilai-Borwein (BB) stepsize iannazzo2018riemannian ; 2013AWen as follows
$l_{k,i}=\begin{cases}\min\{\max\{l_{\min},(\rho_{k}+\gamma_{k})|\frac{\langle\Delta X,\Delta G\rangle}{\|\Delta X\|^{2}}|)\},l_{\max}\},&if~{}\Delta X\neq 0,\\ l_{\max},&if~{}\Delta X=0,\end{cases}$
where $\Delta X=X_{k,i}-X_{k,i-1}$ and $\Delta G=\mathrm{grad}\Phi_{k}(x_{k,i})-\mathrm{grad}\Phi_{k}(x_{k,i-1})$. Moreover, inspired by xuRiemannianAlternatingDescent2024 , for given constant $\theta>1$, $\tau_{1}\in(0,1)$ and $\tau_{2}\in(0,1)$, we set $\rho_{0}=\xi_{0}$ and for $k\geq 1$,
@@
In what follows, we conduct convergence analysis for MPGDA algorithm. First, we present a technical lemma that guarantees the boundedness of the mapping $\bar{y}_{k}$.
@@
Let the mapping $\bar{y}_{k}:\mathbb{E}_{1}\to\mathbb{E}_{2}$ be defined in (15) for $k\in N$. Then for all $x$, $\bar{x}\in\mathbb{E}_{1}$ and $y\in\mathbb{E}_{2}$ , it holds that
@@
$\leq f(x,\bar{y}_{k}(\bar{x}))-g(\bar{y}_{k}(\bar{x}))-\frac{\rho_{k}}{2}\|\bar{y}_{k}(\bar{x})-y_{k}\|^{2}-\frac{\gamma_{k}}{2}\|\bar{y}_{k}(\bar{x})\|^{2}+\frac{\gamma_{k}}{2}\|y\|^{2}$
@@
According to the definition of $\bar{y}_{k}(\bar{x})$, we have
$0\in\nabla_{y}f(\bar{x},\bar{y}_{k}(\bar{x}))-\partial g(\bar{y}_{k}(\bar{x}))-\partial\delta_{S}(\bar{y}_{k}(\bar{x}))-\rho_{k}(\bar{y}_{k}(\bar{x})-y_{k})-\gamma_{k}\bar{y}_{k}(\bar{x}).$
By subgradient inequality, we obtain

### v2

@@ 3 Optimality Conditions
In this section, we introduce two important concepts of stationarity for problem (1), namely, the optimization stationarity and game stationarity, and examine their relationships. To this end, we first define the value function $\Phi:\mathbb{E}_{1}\to\mathbb{R}$ by
@@
The next proposition concerns the value function $\Phi$.
@@
The value function $\Phi$ defined in (9) is $L_{x}$-weakly convex and for all $x\in\mathbb{E}_{1}$, it holds that $\nabla_{x}f(x,y_{x})\in\partial\Phi(x)$ with $y_{x}\in\mathop{\mathrm{argmax}}_{y\in S}\{f(x,y)-g(y)\}$. Moreover, if f(x,y)g(y)f(x,y)-g(y) is strongly concave in $y$, then $\Phi$ is differentiable and $\nabla\Phi(x)=\nabla_{x}f(x,y_{x})$.
@@
Because $f(x,y)$ is $L_{x}$-smooth with respect to $x$, $f(x,y)+\frac{L_{x}}{2}\left\|x\right\|^{2}-g(y)$ is convex in $x$ for any given $y\in S$. Thus, $\Phi(x)+\frac{L_{x}}{2}\left\|x\right\|^{2}=\max_{y\in S}f(x,y)+\frac{L_{x}}{2}\left\|x\right\|^{2}-g(y)$ is convex. Therefore, $\Phi(x)$ is $L_{x}$-weakly convex. Since $S$ is compact, it follows from Danskin’s theorem (bernhardTheoremDanskinApplication1995, Theorem D1) that $\nabla_{x}f(x,y_{x})\in\partial\Phi(x)$, and in particular, $\nabla\Phi(x)=\nabla_{x}f(x,y_{x})$ when f(x,y)g(y)f(x,y)-g(y) is strongly concave in $y$. ∎
In view of Proposition 2 and (1983Optimization, Proposition 2.3.3, Proposition 2.3.6,Proposition 2.4.3), we see that any local minimizer $x^{*}$ of problem (10) must satisfy
@@
This naturally leads to the following definition of optimization stationarity, where the $\varepsilon$-optimization-stationarity is partly motivated by chenProximalGradientMethod2020 and perez2019subdifferential.
@@
    Let $\varepsilon>0$ be given. We say that $x^{*}$ is an $\varepsilon$-optimization-stationary point of problem (1) if there exists $u\in\mathbb{E}$ such that
    $\max\{\|u\|,\mathrm{dist}(0,\partial_{\varepsilon}\Phi(x^{*}+u)+\partial h(x^{*}+u)+\partial\delta_{\mathcal{M}}(x^{*}))\}<\varepsilon,$
    where
    $\partial_{\varepsilon}\Phi(x):=\{z\in\mathbb{E}\mid\Phi(x)+\langle z,y-x\rangle-\varepsilon\leq\Phi(y)+\frac{L_{x}}{2}\|y-x\|^{2},~\forall y\in\mathbb{E}\}.$
@@
Suppose that $x^{*}$ is an optimization-stationary point of problem (1). By Proposition 2, $\Phi(x)+h(x)+L_{x}\|x-x^{*}\|^{2}$ is strongly convex. As $0\in\partial\delta_{\mathcal{M}}(x^{*})+\partial\Phi(x^{*})+\partial h(x^{*})$, and $\partial\delta_{\mathcal{M}}(x^{*})=N_{x^{*}}{\mathcal{M}}=\partial\delta_{T_{x^{*}}{\mathcal{M}}}(0)$, then $x^{*}$ is the unique minimum point of the problem
@@
$f(x^{*},y^{*})+h(x^{*})-g(y^{*})$
$=$ $\inf_{\begin{subarray}{c}x\in A_{x^{*}}{\mathcal{M}},\\ \|x-x^{*}\|\leq 1\end{subarray}}\sup_{y\in S}f(x,y)+h(x)-g(y)+L_{x}\|x-x^{*}\|^{2}$
@@
Conversely, suppose that $(x^{*},y^{*})$ is a game-stationary point of problem (1). Then, according to Proposition 2, $y^{*}=\mathop{\mathrm{argmax}}_{y\in S}f(x^{*},y)-g(y)$ and $\nabla_{x}f(x^{*},y^{*})\in\partial\Phi(x^{*})$. Thus, $0\in\partial\Phi(x^{*})+\partial h(x^{*})+\partial\delta_{\mathcal{M}}(x^{*})$. ∎
@@
where $\beta>0$ and $u=\mathop{\mathrm{argmin}}_{v\in T_{x}\mathcal{M}}\langle\nabla_{x}f(x,y),v\rangle+h(x+v)+\frac{\beta}{2}\|v\|^{2}$. This definition is partly motivated by the stationarity measure used in the Euclidean minimax problem luHybridBlockSuccessive2020. It is straightforward to verify that $(x^{*},y^{*})$ is a game-stationary point of problem (1) if and only if there exists $\beta>0$ such that $\mathcal{G}^{\beta}(x^{*},y^{*})=0$. Using this measure, we can also introduce the notion of $\varepsilon$-game-stationary point.
@@
Let $\varepsilon>0$ be given. Suppose that $(x^{*},y^{*})$ is an $\varepsilon$-game-stationary point of problem (1) with $\mathcal{G}^{\beta}(x^{*},y^{*})\leq\varepsilon$. Then, $x^{*}$ is a $(C\varepsilon)$-optimization-stationary point with
$C=\max\{4\sigma_{y}^{2}+1+2L_{y}\sigma_{y}/\beta,1/\beta,1+L_{x}/\beta\},$
where $L_{x},~L_{y},~\sigma_{y}$ are defined in Assumption 1.
@@
Suppose the point $(x^{*},y^{*})$ is an $\varepsilon$-game-stationary point of problem (1), and let $u^{*}:=\arg\min_{v\in T_{x^{*}}\mathcal{M}}\langle\nabla_{x}f(x^{*},y^{*}),v\rangle+h(x^{*}+v)+\frac{\beta}{2}\|v\|^{2}$. Then, $\|\beta u^{*}\|\leq\varepsilon$, and there exists $z^{*}\in\mathbb{E}_{2}$ such that $\|z^{*}\|\leq 1$ and
$\varepsilon z^{*}\in\nabla_{y}f(x^{*},y^{*})-\partial g(y^{*})-\partial\delta_{S}(y^{*}).$ (14)
We first prove that
$\nabla_{x}f(x^{*}+u^{*},y^{*})\in\partial_{\hat{\varepsilon}}\Phi(x^{*}+u^{*})\text{ with }\hat{\varepsilon}=\varepsilon(4\sigma_{y}^{2}+1)+2L_{y}\sigma_{y}\|u^{*}\|.$ (15)
Consider the auxiliary function
$\Phi_{\varepsilon}^{u^{*}}(x):=$ $\max_{y\in S}f(x+u^{*},y)-g(y)-\frac{\varepsilon}{2}\|y-(y^{*}-z^{*})\|^{2}$
$+\langle\nabla_{y}f(x^{*},y^{*})-\nabla_{y}f(x^{*}+u^{*},y^{*}),y\rangle.$
By (14), the point $y^{*}$ satisfies the first-order optimality condition for the maximization problem defining $\Phi^{u}_{\varepsilon}(x^{*})$. Hence, applying Danskin’s theorem (bernhardTheoremDanskinApplication1995, Theorem D1), we obtain $\nabla\Phi_{\varepsilon}^{u^{*}}(x^{*})=\nabla_{x}f(x^{*}+u^{*},y^{*})$.
Following an argument similar to the proof of Proposition 2, we deduce that $\Phi_{\varepsilon}^{u^{*}}$ is $L_{x}$-weakly convex. Consequently, for any $x\in\mathbb{E}$,
$\Phi_{\varepsilon}^{u^{*}}(x^{*})+\langle\nabla_{x}f(x^{*}+u^{*},y^{*}),x-(x^{*}+u^{*})\rangle$ (16)
$\leq$ $\Phi_{\varepsilon}^{u^{*}}(x-u^{*})+\frac{L_{x}}{2}\|x-(x^{*}+u^{*})\|^{2}.$
Recall that $\sigma_{y}$ is the upper bound of $S$ and $\nabla_{y}f(x,y)$ is $L_{y}$-Lipschitz continuous on $\mathcal{M}\times S$. Using $\|z^{*}\|\leq 1$, we obtain the estimates
$\frac{\varepsilon}{2}\|y-(y^{*}-z^{*})\|^{2}\leq\varepsilon(4\sigma_{y}^{2}+1),$
$\langle\nabla_{y}f(x^{*},y^{*})-\nabla_{y}f(x^{*}+u^{*},y^{*}),y\rangle\leq L_{y}\sigma_{y}\|u^{*}\|.$
From the definitions of $\Phi$ and $\Phi_{\varepsilon}^{u^{*}}$, these inequalities yield
$\Phi(x^{*}+u^{*})\leq\Phi_{\varepsilon}^{u^{*}}(x^{*})+\varepsilon(4\sigma_{y}^{2}+1)+L_{y}\sigma_{y}\|u^{*}\|,$ (17)
and for any $x\in\mathbb{E}$,
$\Phi_{\varepsilon}^{u^{*}}(x-u^{*})\leq\Phi(x)+L_{y}\sigma_{y}\|u^{*}\|.$ (18)
Combining (16), (17) and (18), we obtain that for any $x\in\mathbb{E}$,
$\Phi(x^{*}+u^{*})+\langle\nabla_{x}f(x^{*}+u^{*},y^{*}),x-(x^{*}+u^{*})\rangle-\varepsilon(4\sigma_{y}^{2}+1)-2L_{y}\sigma_{y}\|u^{*}\|$
$\leq$ $\Phi(x)+\frac{L_{x}}{2}\|x-(x^{*}+u^{*})\|^{2}.$
By the definition of the $\varepsilon$-subdifferential $\partial_{\varepsilon}\Phi(\cdot)$, this proves (15).
On the other hand, from the definition of $u^{*}$, we have
$\beta u^{*}-\nabla_{x}f(x^{*},y^{*})\in\partial h(x^{*}+u^{*})+\partial\delta_{\mathcal{M}}(x^{*}).$
Combining this with (15) and recalling that $\nabla_{x}f(\cdot,y)$ is $L_{x}$-Lipschitz continuous, we deduce that
$\mathrm{dist}(0,\partial_{\hat{\varepsilon}}\Phi(x^{*}+u^{*})+\partial h(x^{*}+u^{*})+\partial\delta_{\mathcal{M}}(x^{*}))<(\beta+L_{x})\|u^{*}\|.$
Since $\|\beta u^{*}\|\leq\varepsilon$ by the $\varepsilon$-game-stationarity condition, it follows that $\hat{\varepsilon}\leq\varepsilon(4\sigma_{y}^{2}+1+2L_{y}\sigma_{y}/\beta)$ and $x^{*}$ is a $(C\varepsilon)$-optimization-stationary point with $C=\max\{4\sigma_{y}^{2}+1+2L_{y}\sigma_{y}/\beta,1/\beta,1+L_{x}/\beta\}$. ∎
4 The proposed MPGDA-PA algorithm
In this section, we propose the MPGDA-PA method for solving problem (1) and establish its iteration complexity. Our algorithm development is partially motivated by heApproximationProximalGradient2024; luHybridBlockSuccessive2020; xuRiemannianAlternatingDescent2024. For the $k$-th iteration, we introduce the following value function:
@@ 4 The proposed MPGDA-PA algorithm
where $\Phi_{k}:\mathbb{E}_{1}\to\mathbb{R}$ is defined by
@@
Here, $y_{k}$ denotes the iterate with respect to $y$ obtained in the previous iteration, $\gamma_{k}>0$ is the regularization parameter, and $\rho_{k}>0$ is the proximal parameter. Intuitively, the regularization term $\frac{\gamma_{k}}{2}\|y\|^{2}$ ensures strong concavity of the optimization problem in (9), while the proximal term $\frac{\rho_{k}}{2}\|y-y_{k}\|^{2}$ stabilizes the update of $y$. Moreover, we choose $\gamma_{k}=\gamma_{0}/k^{1/3}$ with some $\gamma_{0}>0$ and assume the non-increasing sequence $\{\rho_{k}\}$ to be summable, i.e.,
@@
We further introduce the mapping $\overline{y}_{k}:\mathbb{E}_{1}\to\mathbb{E}_{2}$ defined by
$\overline{y}_{k}(x):=\mathop{\mathrm{argmax}}_{y\in S}\{f(x,y)-g(y)-\frac{\gamma_{k}}{2}\|y\|^{2}-\frac{\rho_{k}}{2}\|y-y_{k}\|^{2}\}.$ (21)
Then it holds that $\Phi_{k}(x)=f(x,\overline{y}_{k}(x))-\frac{\gamma_{k}}{2}\|\overline{y}_{k}(x)\|^{2}-\frac{\rho_{k}}{2}\|\overline{y}_{k}(x)-y_{k}\|^{2}$ and in view of Danskin’s theorem (danskinTheoryMaxMinApplications1966, Theorem D1), $\Phi_{k}$ is smooth on $\mathbb{E}_{1}$ and $\nabla\Phi_{k}(x)=\nabla_{x}f(x,\overline{y}_{k}(x))$.
Now, at the $k$-th iteration, the proposed MPGDA-PA method computes the new iterate by approximately solving the subproblem:
@@
Inspired by the structure of $Q_{k}$ and the ManPG algorithm (chenProximalGradientMethod2020, Algorithm 4.1), we perform Riemannian manifold proximal gradient descent steps for $T_{k}$ times to generate $x_{k+1}$ with $T_{k}$ being an integer in $1,\overline{T}]$ where $\overline{T}$ is a predetermined positive integer. Then $y_{k+1}$ is updated as $\overline{y}_{k}(x_{k+1})$ with $\overline{y}_{k}$ defined in (21). Note that the update of $y$ can also be interpreted as performing a proximal ascent step for the function $f(x_{k+1},y)-g(y)-\frac{\gamma_{k}}{2}\|y\|^{2}$ on $S$. We formally present the proposed method in Algorithm 1.
Algorithm 1 Manifold Proximal Gradient Descent Ascent (MPGDA-PA) for problem (1)
[…]
1:for $k=0,1,\dots$ do
2:  Set $x_{k,0}=x_{k}$.
@@
$v_{k,i}:=\mathop{\mathrm{argmin}}_{v\in T_{x_{k,i}}\mathcal{M}}\langle\nabla_{x}f(x_{k,i},\overline{y}_{k}(x_{k,i})),v\rangle+h(x_{k,i}+v)+\frac{\beta_{k,i}}{2}\left\|v\right\|^{2},$ (23)
@@
Before proceeding, we make some remarks on the iteration procedure of the MPGDA-PA algorithm. First, given $x_{k,i}\in\mathcal{M}$, the MPGDA-PA algorithm computes a descent direction $v_{k,i}$ of $Q_{k}$ restricted to the tangent space $T_{x_{k,i}}\mathcal{M}$ via tackling the linear constrained strongly convex optimization problem (23), which can usually be efficiently solved, see (li2024proximal, Section 4), for example. Second, for all $x\in\mathcal{M}$, $\overline{y}_{k}(x)$ defined in (21) can be solved by applying FISTA Beck2017First to problem (20), which converges linearly due to the strong convexity and smoothness of
$\frac{\gamma_{k}}{2}\|\cdot\|^{2}+\frac{\rho_{k}}{2}\|\cdot-y_{k}\|^{2}-f(x,\cdot).$
In particular, if $f(x,y)$ is linear with respect to $y$ for fixed $x$, i.e., $f(x,y)=\langle\mathcal{A}(x),y\rangle$, then $\overline{y}_{k}(x)$ can be further simplified into a proximal problem like (2b) as follows:
$\overline{y}_{k}(x)=\mathop{\mathrm{argmax}}_{y\in S}\left\{-g(y)-\frac{\gamma_{k}+\rho_{k}}{2}\|y-\frac{\rho_{k}y_{k}+\mathcal{A}(x)}{\rho_{k}+\gamma_{k}}\|^{2}\right\}.$
Third, in practice, we make some choices of parameters $l_{k,i}$ and $\rho_{k}$, which enhance the efficiency of the MPGDA-PA algorithm in our test and also satisfy $l_{k,i}\in[l_{\min},l_{\max}]$ and $\rho_{k}\in S$. Motivated by the Riemannian Barzilai-Borwein (BB) stepsize iannazzo2018riemannian; 2013AWen, we choose $l_{k,i}$ as follows:
$l_{k,i}=\begin{cases}\min\{\max\{l_{\min},(\rho_{k}+\gamma_{k})|\frac{\langle\Delta X,\Delta R\rangle}{\|\Delta X\|^{2}}|)\},l_{\max}\},&if~\Delta X\neq 0,\\ l_{\max},&if~\Delta X=0,\end{cases}$
where $\Delta X=X_{k,i}-X_{k,i-1}$ and $\Delta R=\mathrm{grad}\Phi_{k}(x_{k,i})-\mathrm{grad}\Phi_{k}(x_{k,i-1})$. Moreover, inspired by xuRiemannianAlternatingDescent2024, for given constant $\theta>1$, $\tau_{1}\in(0,1)$ and $\tau_{2}\in(0,1)$, we set $\rho_{0}=\xi_{0}$ and for $k\geq 1$,
@@
In what follows, we conduct convergence analysis for the MPGDA-PA algorithm. First, we present a technical lemma that guarantees the boundedness of the mapping $\overline{y}_{k}$.
@@
Let the mapping $\overline{y}_{k}:\mathbb{E}_{1}\to\mathbb{E}_{2}$ be defined in (21) for $k\in\mathbb{N}$. Then for all $x$, $\bar{x}\in\mathbb{E}_{1}$ and $y\in\mathbb{E}_{2}$ , it holds that
@@
$\leq f(x,\overline{y}_{k}(\bar{x}))-g(\overline{y}_{k}(\bar{x}))-\frac{\rho_{k}}{2}\|\overline{y}_{k}(\bar{x})-y_{k}\|^{2}-\frac{\gamma_{k}}{2}\|\overline{y}_{k}(\bar{x})\|^{2}+\frac{\gamma_{k}}{2}\|y\|^{2}$
@@
By the definition of $\overline{y}_{k}(\bar{x})$, we have
$0\in\nabla_{y}f(\bar{x},\overline{y}_{k}(\bar{x}))-\partial g(\overline{y}_{k}(\bar{x}))-\partial\delta_{S}(\overline{y}_{k}(\bar{x}))-\rho_{k}(\overline{y}_{k}(\bar{x})-y_{k})-\gamma_{k}\overline{y}_{k}(\bar{x}).$
This and the concavity of $f(x,\cdot)-g(\cdot)$ yield that